An actor invests some money at 7%, and $41000and more than four times the amount of 10%. The total annual interest earned from the investment is $44990. How much did he invest at each amount? Use the six step method.
He invested ? at 7%
He invested ? at 10%

Answer by richwmiller(17219)

Total yearly interest for the two accounts is: $44990
x=4y+41000
0.07*x+0.1*y=44990
Substitute for x
0.07*(4y+41000)+0.1*y=44990
Multiply out
0.28y+2870+0.1*y=44990
Combine like terms.
0.28y+2870+0.1*y=44990
0.28y+0.1*y=44990-2870
0.38*y=42120
Isolate y
0.38*y=42120
y=42120/0.38
y=$110842.105 at 10%
rounded off
y=$110842.11 at 10%
x=4y+41000
x=4*110842.105+41000
x=443368.421+41000
x=$484368.421 at 7%
rounded off
x=$484368.42 at 7%